# Optimally-Weighted Maximum Mean Discrepancy Framework for Continual Learning

### v1

@@ Abstract
Continual learning has emerged as a pivotal area of research, primarily due to its advantageous characteristic that allows models to persistently acquire and retain information. However, catastrophic forgetting can severely impair model performance. In this study, we tackle the issue of network forgetting by introducing a novel framework termed Optimally-Weighted Maximum Mean Discrepancy (OWMMD), which imposes penalties on representation alterations via a Multi-Level Feature Matching Mechanism (MLFMM). Furthermore, we propose an Adaptive Regularization Optimization (ARO) strategy to refine the adaptive weight vectors, which autonomously assess the significance of each feature layer throughout the optimization process. We conduct a comprehensive series of experiments, benchmarking our proposed method against several established baselines. The empirical findings indicate that our approach achieves state-of-the-art performance.
@@ 1 Introduction
Rehearsal-based approaches retain a subset of historical data and utilize it during the acquisition of new tasks [3, 61]. While these strategies facilitate the retention of knowledge from previous tasks, they frequently encounter constraints related to memory capacity and computational inefficiencies. In contrast, regularization-based techniques [22] seek to mitigate substantial alterations to essential model parameters by implementing constraints that penalize significant changes. Approaches such as Elastic Weight Consolidation (EWC) [22] and Synaptic Intelligence (SI) [65] exemplify this category, demonstrating promising outcomes by safeguarding critical weights across various tasks. Nevertheless, these methodologies predominantly emphasize weight preservation and may face challenges in addressing the heterogeneous and dynamic nature of data encountered in practical applications.
Recent developments in Mixture-of-Experts (MoE) models and attention mechanisms [60] have showcased their efficacy in tackling continual learning (CL) challenges. These methodologies enable the model to allocate computational resources in a targeted manner, thereby enhancing both efficiency and performance in multi-task scenarios. However, despite these innovations, current solutions frequently encounter trade-offs related to complexity, memory consumption, and the ability to generalize across diverse tasks.
@@
Recent advancements in continual learning have increasingly emphasized knowledge distillation-based methodologies [17, 20, 40], wherein the fundamental principle involves transferring knowledge from a previously trained model to the model currently being trained on the active task. In this context, the knowledge distillation framework typically comprises a teacher module, which is trained on the prior task and subsequently frozen, alongside a student module that undergoes continuous training on the current task. A regularization term is incorporated to minimize the divergence between the outputs of the teacher and student, thereby regulating the model’s optimization process to mitigate the risk of catastrophic forgetting. The efficacy of knowledge distillation methods is underscored by their capacity to maintain performance on earlier tasks without necessitating the retention of any memorized samples, establishing them as a prominent strategy for addressing challenges in continual learning. Nonetheless, a significant limitation of knowledge distillation approaches is their propensity to impair the learning capacity for new tasks, attributable to the issue of over-regularization.
@@
In this paper, we tackle network forgetting in continual learning by proposing an innovative framework known as Optimally-Weighted Maximum Mean Discrepancy (OWMMD). This framework is designed to mitigate representation shifts through a probability-based distance measure. We introduce a novel Multi-Level Feature Matching Mechanism (MLFMM) that aims to minimize the distance between previously and currently learned representations throughout the model’s optimization process, effectively addressing the challenge of network forgetting. Unlike conventional distillation techniques that concentrate on aligning final outputs, MLFMM penalizes alterations in representations across all multi-level feature layers, leading to enhanced performance. To address the problem of over-regularization, MLFMM incorporates a new Adaptive Regularization Optimization (ARO) strategy that automatically assesses the significance of each feature layer during the regularization phase, thereby preventing network forgetting while facilitating the learning of future tasks. We conduct a comprehensive series of experiments and benchmark our method against several contemporary baselines. The empirical findings indicate that our proposed approach achieves state-of-the-art performance.
@@
    We introduce an Adaptive Regularization Optimization (ARO) framework that enables the model to evaluate the significance of each feature layer in real-time throughout the optimization process, thereby guaranteeing optimal regularization.
@@
    We perform comprehensive experiments to assess the efficacy of our methodology, benchmarking it against multiple recognized baselines in continual learning. The findings indicate that our approach attains state-of-the-art performance, successfully alleviating the issue of forgetting while preserving high accuracy.
The subsequent sections of this manuscript are organized as follows: Chapter 2 provides an overview of the Related Work in continual learning, critically evaluating the advantages and drawbacks of current methodologies. Chapter 3 outlines the Methodology, elaborating on the MLFMM framework and the adaptive regularization optimization component. In Chapter 4, we detail the experimental design and outcomes, showcasing the efficacy of our approach across multiple continual learning benchmarks. Lastly, Chapter 5 wraps up the paper with a synthesis of the findings and suggestions for future research avenues.
@@ 2 Related Work
Various methods have been proposed to tackle the challenge of catastrophic forgetting, often categorizing into approaches such as rehearsal-based methods, knowledge distillation, regularization-based methods, and architecture-based methods.
Rehearsal-based methods store a small subset of past data, often referred to as an episodic memory, and replay this data during training on new tasks to retain knowledge from previous tasks [4, 9, 32]. ER [10] demonstrates that even very small memory buffers can significantly improve generalization. The study found that training with just one example per class from previous tasks led to improvements about 10% performance across various benchmarks, suggesting that simple memory-based methods can outperform more complex continual learning algorithms. GEM [32] mitigates forgetting by constraining the optimization process to avoid interference with previous tasks. GEM allows for beneficial transfer of knowledge to new tasks, showing strong performance on datasets such as MNIST [26, 37] and CIFAR-100 [45]. A-GEM [9] builds on GEM but improves its computational and memory efficiency, providing a better trade-off between performance and resource consumption. A-GEM also demonstrated the ability to learn more efficiently when provided with task descriptors. GSS [2] proposes an efficient way to select samples for replay buffers by maximizing the diversity of samples. This method frames sample selection as a constraint reduction problem, where the goal is to choose a fixed subset of constraints that best approximate the feasible region defined by the original constraints. This approach outperforms traditional task-boundary-based methods in terms of accuracy and efficiency. HAL [8] introduces a bilevel optimization technique that complements experience replay by anchoring the knowledge from past tasks. By preserving predictions on anchor points through fine-tuning on episodic memory, this method improves both accuracy and the mitigation of forgetting compared to standard experience replay.
@@
Regularization-based methods seek to prevent catastrophic forgetting by restricting how much the model’s parameters can change during training [24, 59]. EWC [22] presents a seminal approach for alleviating forgetting in neural networks. The method involves selective slowing down of learning on weights that are important for previous tasks. By preserving the important weights through a penalty on large weight changes, the model can continue learning new tasks while maintaining knowledge from previous ones. This regularization approach was demonstrated on both classification tasks (MNIST) and sequential learning tasks (Atari 2600 games), showing that the model can successfully retain knowledge even after long periods without encountering the original tasks. oEWC [43] introduces a regularization strategy within a progress and compress framework. The method works by partitioning the network into a knowledge base (which stores information about previous tasks) and an active column (which focuses on the current task). After a new task is learned, the active column is consolidated into the knowledge base, protecting the knowledge acquired so far. The key aspect of this approach is its ability to achieve this consolidation without growing the architecture, storing old data, or requiring task-specific parameters. It is shown to work effectively on sequential classification tasks and reinforcement learning domains like Atari games and maze navigation, providing a balance between learning new tasks and preserving old knowledge. SI [65] mimics the adaptability of biological neural networks. In this approach, each synapse accumulates task-relevant information over time, allowing the model to store new memories while maintaining knowledge from prior tasks. The accumulated task-specific information helps the model regulate how much each synapse can change, ensuring that learning of new tasks does not interfere with older knowledge. SI significantly reduces forgetting and provides a computationally efficient method for continual learning, as demonstrated in classification tasks on benchmark datasets like MNIST. RW incorporates a KL-divergence-based perspective to measure the difference between the current model and the previous task’s knowledge. The method introduces two new metrics—forgetting and intransigence—to quantify how well an algorithm balances retaining old knowledge and updating for new tasks. The results show that RW offers superior performance compared to traditional methods, as it provides a better trade-off between forgetting and intransigence. The approach is evaluated on several benchmark datasets (MNIST, CIFAR-100), with RW showing significant improvement in preserving knowledge while being able to update for new tasks.
@@
Architecture-based approaches tackle catastrophic forgetting by altering the network structure to accommodate new tasks [61, 13, 63, 18, 64, 1]. Progressive Neural Networks (PNN) [41] incrementally adds new modules to the network for each task, leading to a growing memory requirement. PackNet [33] introduces a method that allows adding multiple tasks to a single deep network by exploiting redundancies in large models through iterative pruning. This approach works by pruning a part of the network to free up parameters, which are then used for learning new tasks. The key feature of PackNet is its ability to optimize for the current task without sacrificing performance on previous tasks. Extensive experiments, such as adding multiple fine-grained classification tasks to a single ImageNet-trained VGG-16 network, demonstrate that this method provides better robustness against catastrophic forgetting compared to previous approaches that relied on proxy losses. HAT [44] proposes a task-based hard attention mechanism to prevent catastrophic forgetting. In this approach, a hard attention mask is learned for each task during training, allowing the network to focus on the relevant parameters for each specific task while ignoring irrelevant ones. This prevents forgetting by restricting updates to previously learned tasks. The hard attention mask is updated during training, with previous masks conditioning the learning of new tasks. This method significantly reduces forgetting and provides flexibility in controlling the stability and compactness of learned knowledge, making it suitable for both online learning and network compression applications. Quang Pham et al [38] introduce a DualNets framework based on the Complementary Learning Systems (CLS) theory from neuroscience, which posits that human learning occurs through two complementary systems: a fast learning system for individual experiences (hippocampus) and a slow learning system for structured knowledge (neocortex). DualNets implements this idea in deep neural networks by dividing the network into two components: a fast learning system for supervised learning of task-specific representations and a slow learning system for learning task-agnostic, general representations through Self-Supervised Learning (SSL). This dual approach helps the model balance learning efficiency with retention of previously learned knowledge, making it effective in both task-aware and task-free continual learning scenarios. DualNets has shown strong performance on benchmarks like CTrL, outperforming dynamic architecture methods in some cases. Hongbo et al explore the MoE [28] architecture in the context of continual learning, providing the first theoretical analysis of MoE’s impact in this domain. MoE models use a collection of specialized experts, with a router selecting the most appropriate expert for each task. The paper shows that MoE can diversify its experts to handle different tasks and balance the workload across experts. The study suggests that MoE in continual learning may require termination of updates to the gating network after sufficient training rounds for system convergence, a condition that is not required in non-continual MoE studies. Additionally, the paper provides insights into expected forgetting and generalization error in MoE, highlighting that adding more experts can delay convergence without improving performance. The theoretical insights are validated through experiments on both synthetic and real-world datasets, demonstrating the potential benefits of MoE in continual learning for deep neural networks (DNNs).
@@
Rehearsal-based Store and replay past data to retain knowledge from previous tasks. ER [10], GEM [32], A-GEM [9], GSS [2], HAL [8]
Knowledge Distillation Transfers knowledge from a teacher model to a student model to prevent forgetting. LWF [31], iCaRL [39], DER [34]
Regularization-based Restricts changes to model parameters to avoid catastrophic forgetting. EWC [22], oEWC [43], SI [65], RW [7]
Architecture-based Alters network structure to accommodate new tasks. PNN [41], PackNet [33], HAT [44], DualNets [38], MoE [28]

### v4

# Learning Dynamic Representations via An Optimally-Weighted Maximum Mean Discrepancy Optimization Framework for Continual Learning

@@ Abstract
Continual learning has emerged as a pivotal area of research, primarily due to its advantageous characteristic that allows models to persistently acquire and retain information. However, catastrophic forgetting can severely impair model performance. In this study, we address network forgetting by introducing a novel framework termed Optimally-Weighted Maximum Mean Discrepancy (OWMMD), which imposes penalties on representation alterations via a Multi-Level Feature Matching Mechanism (MLFMM). Furthermore, we propose an Adaptive Regularization Optimization (ARO) strategy to refine the adaptive weight vectors, which autonomously assess the significance of each feature layer throughout the optimization process, The proposed ARO approach can relieve the over-regularization problem and promote the future task learning. We conduct a comprehensive series of experiments, benchmarking our proposed method against several established baselines. The empirical findings indicate that our approach achieves state-of-the-art performance.
keywords:
Continual Learning , Multi-Level Feature Matching , Adaptive Regularization Optimization
journal: Knowledge-Based Systems
[…]
{highlights}
An innovative framework termed Optimally Weighted Maximum Mean Discrepancy (OWMMD) is proposed to mitigate catastrophic forgetting in continual learning paradigms.
A Multi-Level Feature Matching Mechanism (MLFMM) is pro impose penalties on the modification of feature representations across various tasks
An Adaptive Regularization Optimization (ARO) framework that enables the model to evaluate the significance of each feature layer in real-time throughout the optimization process.
@@ 1 Introduction
Recent research has introduced various methodologies to tackle the issue of network forgetting in continual learning, which can be categorized into three primary approaches: memory-based methods [14, 15, 16], dynamic expansion-based methods [17, 18], and regularization-based methods [19, 20]. Among these, the rehearsal-based strategy is a straightforward yet effective technique to mitigate network forgetting, focusing on preserving a subset of historical data. These stored samples are reutilized and integrated with new data to optimize the model during the learning of new tasks [21, 22]. However, the efficacy of rehearsal-based methods is heavily dependent on the quality of the retained samples. Regularization-based techniques [23, 24] aim to prevent significant changes to crucial model parameters by incorporating additional regularization terms into the main objective function. Despite this, regularization-based methods primarily focus on weight conservation and may struggle with the diverse and evolving nature of data encountered in real-world scenarios. Recent advancements have proposed employing dynamic expansion models to address network forgetting in continual learning [5]. These methods dynamically generate new sub-networks and hidden layers to assimilate new information over time while preserving all previously trained parameters to ensure optimal performance on earlier tasks. Nonetheless, a notable limitation of the dynamic expansion model is the increasing complexity of the model as the number of tasks expands.
@@
Recent advancements in continual learning have increasingly emphasized knowledge distillation-based methodologies [25, 26, 27], wherein the fundamental principle involves transferring knowledge from a previously trained model to the model currently being trained on the active task. In this context, the knowledge distillation framework typically comprises a teacher module, which is trained on the prior task and subsequently frozen, alongside a student module that undergoes continuous training on the current task. A regularization term is incorporated to minimize the divergence between the outputs of the teacher and student, thereby regulating the model’s optimization process to mitigate the risk of catastrophic forgetting. The efficacy of knowledge distillation methods is demonstrated by their capacity to maintain performance on earlier tasks without necessitating the retention of any memorized samples, establishing them as a prominent strategy for addressing challenges in continual learning. Nonetheless, a significant limitation of knowledge distillation approaches is their propensity to impair the learning capacity for new tasks, leading to the over-regularization problem.
@@
In this paper, we tackle network forgetting in continual learning by proposing an innovative framework known as Optimally-Weighted Maximum Mean Discrepancy (OWMMD). This framework is designed to mitigate representation shifts through a probability-based distance measure. Specifically, a novel Multi-Level Feature Matching Mechanism (MLFMM) is proposed to minimize the distance between previously and currently learned representations throughout the model’s optimization process, effectively addressing the challenge of network forgetting. Unlike conventional distillation techniques that concentrate on aligning final outputs, MLFMM penalizes alterations in representations across all multi-level feature layers, which can further relieve network forgetting. To address the problem of over-regularization, OWMMD incorporates a new Adaptive Regularization Optimization (ARO) strategy that automatically assesses the significance of each feature layer during the model optimization process, thereby preventing network forgetting while facilitating the learning of future tasks. We conduct a comprehensive series of experiments and benchmark our method against several contemporary baselines. The empirical findings indicate that our proposed approach achieves state-of-the-art performance.
@@
    We introduce an Adaptive Regularization Optimization (ARO) framework that enables the model to evaluate the significance of each feature layer in real-time throughout the optimization process, thereby guaranteeing an optimal regularization process.
@@
    We perform comprehensive experiments to assess the efficacy of our methodology, benchmarking it against multiple recognized baselines in continual learning. The findings indicate that our approach achieves state-of-the-art performance, successfully alleviating the problem of forgetting while preserving high accuracy.
The subsequent sections of this manuscript are organized as follows: Chapter 2 provides an overview of the Related Work in continual learning, critically evaluating the advantages and drawbacks of current methodologies. Chapter 3 outlines the Methodology, elaborating on the OWMMD framework and the adaptive regularization optimization component. In Chapter 4, we detail the experimental design and outcomes, showcasing the efficacy of our approach across multiple continual learning benchmarks. Lastly, Chapter 5 wraps up the paper with a synthesis of the findings and suggestions for future research avenues.
@@ 2 Related Work
Various methods have been proposed to tackle the challenge of catastrophic forgetting, often categorizing into approaches such as rehearsal-based methods, knowledge distillation, regularization-based methods, and architecture-based methods. We summarize many important baselines in Table 1.
Rehearsal-based methods store a small subset of past data, often referred to as an episodic memory, and replay this data during training on new tasks to retain knowledge from previous tasks [14, 15, 16]. ER [28] demonstrates that even very small memory buffers can significantly improve generalization. The study found that training with just one example per class from previous tasks led to improvements about 10% performance across various benchmarks, suggesting that simple memory-based methods can outperform more complex continual learning algorithms. GEM [16] mitigates forgetting by constraining the optimization process to avoid interference with previous tasks. GEM allows for beneficial transfer of knowledge to new tasks, showing strong performance on datasets such as MNIST [29, 30] and CIFAR-100 [31]. A-GEM [15] builds on GEM but improves its computational and memory efficiency, providing a better trade-off between performance and resource consumption. A-GEM also demonstrated the ability to learn more efficiently when provided with task descriptors. GSS [32] proposes an efficient way to select samples for replay buffers by maximizing the diversity of samples. This method frames sample selection as a constraint reduction problem, where the goal is to choose a fixed subset of constraints that best approximate the feasible region defined by the original constraints. This approach outperforms traditional task-boundary-based methods in terms of accuracy and efficiency. HAL [33] introduces a bilevel optimization technique that complements experience replay by anchoring the knowledge from past tasks. By preserving predictions on anchor points through fine-tuning on episodic memory, this method improves both accuracy and the mitigation of forgetting compared to standard experience replay. For industrial applications, TRINA [34] introduces a federated continual learning framework with self-challenge rehearsal, which generates historical distributions through masked recovery tasks using random scales and positions. This method enhances the model’s ability to recall complex data distributions while maintaining training stability in industrial monitoring scenarios with spatiotemporal heterogeneity.
@@
Regularization-based methods seek to prevent catastrophic forgetting by restricting how much the model’s parameters can change during training [19, 20]. EWC [23] presents a seminal approach for alleviating forgetting in neural networks. The method involves selective slowing down of learning on weights that are important for previous tasks. By preserving the important weights through a penalty on large weight changes, the model can continue learning new tasks while maintaining knowledge from previous ones. This regularization approach was demonstrated on both classification tasks (MNIST) and sequential learning tasks (Atari 2600 games), showing that the model can successfully retain knowledge even after long periods without encountering the original tasks. oEWC [40] introduces a regularization strategy within a progress and compress framework. The method works by partitioning the network into a knowledge base (which stores information about previous tasks) and an active column (which focuses on the current task). After a new task is learned, the active column is consolidated into the knowledge base, protecting the knowledge acquired so far. The key aspect of this approach is its ability to achieve this consolidation without growing the architecture, storing old data, or requiring task-specific parameters. It is shown to work effectively on sequential classification tasks and reinforcement learning domains like Atari games and maze navigation, providing a balance between learning new tasks and preserving old knowledge. SI [24] mimics the adaptability of biological neural networks. In this approach, each synapse accumulates task-relevant information over time, allowing the model to store new memories while maintaining knowledge from prior tasks. The accumulated task-specific information helps the model regulate how much each synapse can change, ensuring that learning of new tasks does not interfere with older knowledge. SI significantly reduces forgetting and provides a computationally efficient method for continual learning, as demonstrated in classification tasks on benchmark datasets like MNIST. RW [41] incorporates a KL-divergence-based perspective to measure the difference between the current model and the previous task’s knowledge. The method introduces two new metrics, forgetting and intransigence, to quantify how well an algorithm balances retaining old knowledge and updating for new tasks. The results show that RW offers superior performance compared to traditional methods, as it provides a better trade-off between forgetting and intransigence. The approach is evaluated on several benchmark datasets (MNIST, CIFAR-100), with RW showing significant improvement in preserving knowledge while being able to update for new tasks.
@@
Architecture-based approaches tackle catastrophic forgetting by altering the network structure to accommodate new tasks [17, 18, 22, 42, 43, 44]. HAT [45] proposes a task-based hard attention mechanism to prevent catastrophic forgetting. In this approach, a hard attention mask is learned for each task during training, allowing the network to focus on the relevant parameters for each specific task while ignoring irrelevant ones. This prevents forgetting by restricting updates to previously learned tasks. The hard attention mask is updated during training, with previous masks conditioning the learning of new tasks. This method significantly reduces forgetting and provides flexibility in controlling the stability and compactness of learned knowledge, making it suitable for both online learning and network compression applications. Quang Pham et al [46] introduce a DualNets framework based on the Complementary Learning Systems (CLS) theory from neuroscience, which posits that human learning occurs through two complementary systems: a fast learning system for individual experiences (hippocampus) and a slow learning system for structured knowledge (neocortex). DualNets implements this idea in deep neural networks by dividing the network into two components: a fast learning system for supervised learning of task-specific representations and a slow learning system for learning task-agnostic, general representations through Self-Supervised Learning (SSL). This dual approach helps the model balance learning efficiency with retention of previously learned knowledge, making it effective in both task-aware and task-free continual learning scenarios. DualNets has shown strong performance on benchmarks like CTrL, outperforming dynamic architecture methods in some cases. Hongbo et al explore the MoE [47] architecture in the context of continual learning, providing the first theoretical analysis of MoE’s impact in this domain. MoE models use a collection of specialized experts, with a router selecting the most appropriate expert for each task. The paper shows that MoE can diversify its experts to handle different tasks and balance the workload across experts. The study suggests that MoE in continual learning may require termination of updates to the gating network after sufficient training rounds for system convergence, a condition that is not required in non-continual MoE studies. Additionally, the paper provides insights into expected forgetting and generalization error in MoE, highlighting that adding more experts can delay convergence without improving performance. The theoretical insights are validated through experiments on both synthetic and real-world datasets, demonstrating the potential benefits of MoE in continual learning for deep neural networks (DNNs). DBLF [48] constructs dedicated branch layers for old tasks and dynamically fuses them with new task layers via a two-stage training process (adaptation and fusion), effectively addressing model growth in imbalanced rotating machinery fault diagnosis. Lu et al [49] introduce knowledge-guided prompt alignment with contrastive hard negatives, improving task-specific prompt distinguishability through a semantic-enhanced module. For universal information extraction, Jin et al designed a multi-LoRA architecture [50] that freezes pretrained weights and trains independent low-rank adapters (LoRA) for auxiliary and target tasks, merging parameters during inference. WKNN-CLCMTVD [51] continuously updates bio-inspired memory cells using k-nearest neighbor rules, enabling fast adaptation to time-varying data spaces. Svoboda et al [52] combines Hoeffding trees with PELT-based change point detection to dynamically select model ensembles for non-stationary natural gas consumption forecasting.
@@
Rehearsal-based Store and replay past data to retain knowledge from previous tasks. ER [28], GEM [16], A-GEM [15], GSS [32], HAL [33], TRINA [34]
Knowledge Distillation Transfer knowledge from a teacher model to a student model to prevent forgetting. LWF [37], iCaRL [38], DER [53]
Regularization-based Restrict changes to model parameters to avoid catastrophic forgetting. EWC [23], oEWC [40], SI [24], RW [41]
Architecture-based Alter network structure to accommodate new tasks. HAT [45], DualNets [46], MoE [47], DBLF [48], knowledge-guided prompt [49], multi-LoRA [50], WKNN-CLCMTVD [51], [52]
@@ 3 Methodology
3.1 Problem definition
The description of mathematical notations of this paper is showed in Table 2. In continual learning, a model is unable to access the complete training dataset simultaneously and instead acquires knowledge from a continuously evolving data stream. A common scenario in continual learning is referred to as class-incremental learning, which encompasses a sequence of $N$ tasks $\{T_{1},T_{2},\dots,T_{N}\}$. Each task $T_{i}$ is linked to a labeled training dataset $D^{s}_{i}=\{({\bf x}^{i}_{j},{\bf y}^{i}_{j})\,|\,j=1,\cdots,N^{s}_{i}\}$, where ${\bf x}^{i}_{j}\subset\mathcal{X}$ and ${\bf y}^{i}_{j}\subset\mathcal{Y}$ represent the $j$-th paired sample from the $i$-th task, with $N^{s}_{i}$ indicating the total number of samples in the training dataset $D^{s}_{i}$. Here, $\mathcal{X}$ and $\mathcal{Y}$ denote the respective spaces of data samples and labels. Additionally, let $D^{t}_{i}=\{({\bf x}^{i}_{j},{\bf y}^{i}_{j})\,|\,j=1,\cdots,N^{t}_{i}\}$ represent the testing dataset for the $i$-th task, where $N^{t}_{i}$ signifies the total number of testing samples associated with the $i$-th task.
In a continual learning framework, the goal of the model is to identify the optimal parameter set ${\theta^{\star}}$ from the parameter space $\tilde{\Theta}$, which minimizes the training loss across all tasks ${T_{1},\cdots,T_{i}}$ while learning the current task $T_{i}$, as expressed by the following equation :